At 327 K and 1.09 x 10-2 atm, the density of a gas is 1.66 x 10-5 g/cm3. (a) Find vrms for the gas molecules. (b) Find the molar mass of the gas.
Answer #1

a) Vrms = square root (3*P/density) = square root( 3* 1.09*101325*10-2 / 1.66*10-2 ) = 446.76 m/s

b) Vrms = square root ( 3* R*T/M) = square root (3*8.314*327/M)

446.76 = square root (3*8.314*327/M)

on solving, we get M = .04086 kg = 40.86 g = 41 g (approximately)